OpenStudy (anonymous):

HELP ME PLEASEEE. Justin's gas tank is 1/6 full. After he buys 11 gallons of gas, it is 2/3 full. How many gallons can Justin's tank hold?

OpenStudy (anonymous):

\[(1/6)*x(gal)+11(gal)=(2/3)*x(gal) =====> X= 22 Gallons The answer is that Justin's Tanks can hold 22 Gallons of gasoline

OpenStudy (anonymous):

I got 22 also

OpenStudy (anonymous):

The initial "\" is not supposed to be in the equation. Please leave me a medal!

OpenStudy (anonymous):

THANK YOU ! can you help me with another one?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

I guess @danielfife beat me to it. here is the same answer :) http://www.demonstranda.com/book/questionDetails/324